### v1

@@ 5.4 Tier 3: Semantic Level
Time Precision (TP) and Duration Precision (DP) are defined analogously.
@@
Time Average Shift (TAS) and Duration Average Shift (DAS) are defined analogously. Signedness is kept in order to identify the direction in which the underlying OMR system tends to move the notes. These should be accompanied by the corresponding standard deviation.
@@ 6 The COMREF Dataset
We have developed a dataset using our notation built on transcriptions of public domain works as a proof of concept of the notation format. In particular, we have used the OpenScore project’s transcriptions of widely known works such as The Art of the Fugue by J.S. Bach or the Planets by Gustav Holst, among others. We have also incorporated the Lieder Corpus Gotham and Jonas (2022) and the String Quartet corpus str (2023). All these scores are engraved in MusicXML, but we plan on supporting MEI as a conversion source at some point in the future.
@@ 6.1 Images
In summary, the dataset is produced by processing of 896 individual works into images at the measure level (including all staves that belong to it), to produce a total of $439.972$ images. After verification, $3.260$ images are removed due to noise in the annotations, totalling $436.752$ images after cleanup.
@@
The process through which the dataset was produced is summarised in Figure 4. Scores are engraved through Verovio into page-level SVG files. Using the hierarchical structure within the SVG and exploiting the optional identifier information Verovio can be instructed to attach, we employ a splitter script that finds bounding boxes for every measure and engraves them individually. It also finds what measures lie at the beginning of a line in order to properly add this information on the notation. The bounding boxes for all measures are cropped at the next staff vertically and a fixed amount of pixels horizontally.
@@ 6.2 Ground Truth
The dataset incorporates 91 different classes, which are shown alongside their frequencies in Table 3. A relevant aspect of this dataset is that it is quite unbalanced, with roughly 70% of the elements pertaining to the 4 first classes.
Table 3: Tokens and their frequencies in the COMREF dataset.
Token Count Token Count Token Count
notehead_black 1986948 rest_breve 18103 notehead_white_cue 738
beam 1517194 clef_C 17664 rest_whole 681
stem_down 1018045 notehead_black_grace 13086 regular 529
stem_up 803884 tenuto 11444 sfz 462
flag 244232 strong_accent 9414 rf 173
accidental_sharp 226428 pp 8484 mordent 136
accidental_flat 211192 rest_32nd 8381 ffff 130
notehead_white 197408 ff 7449 sfp 129
staccato 171070 sf 7282 notehead_breve 125
slur_stop 168855 light_light 7230 sffz 105
slur_start 168853 notehead_black_cue 6889 clef_G_oct1 102
dot 153128 rest_maxima 6171 turn 83
rest_eighth 126179 mf 5748 caesura 75
tied_start 101092 fermata 5330 slide_start 72
tied_stop 101029 light_heavy 3294 slide_stop 71
accidental_natural 99920 rest_64th 3248 accidental_natural_sharp 51
rest_quarter 94522 trill_mark 2589 rfz 43
clef_G 75140 timesig_common 2366 segno 31
clef_F 46386 accidental_double_sharp 1990 pppp 15
tuplet_stop 43781 fz 1700 glissando_stop 14
tuplet_start 43650 clef_G_oct-1 1295 glissando_start 10
rest_long 43610 staccatissimo 1268 clef_F_oct1 9
wedge_stop 38286 mp 1245 notehead_black_grace_cue 8
accent 27450 wavy_line 1107 coda 7
arpeggiate 22102 fff 987 schleifer 4
rest_half 21554 accidental_flat_flat 969 sfpp 3
rest_16th 21452 fp 913 accidental_natural_flat 2
p 21103 timesig_cut 890 pppppp 1
wedge_crescendo 20115 ppp 862 ppppp 1
f 19771 clef_percussion 774
wedge_diminuendo 18233 clef_F_oct-1 760 Total 8015354
@@ 7 Conclusions
In this paper we have argued for the implementation of a Common Optical Music Recognition Evaluation Framework through the instauration of a notation format in which score reconstruction can be evaluated fairly an unambiguously. Our proposed reification of this idea is the MTN format, which draws heavy inspiration from existing formats and adapts many ideas already present in the field while also being quite expressive and general. Its main idea is exploiting the expresiveness of graphs while keeping the ordering properties of sequences, thanks to which evaluation and inference are better-defined and less computationally demanding. Of course, CWMN is a tremendously complex notation system which has been optimised and streamlined for hundreds of years. It is highly likely that MTN (or any other computational representation of music for that matter) cannot model every possible music score that has been written. Nevertheless, we believe the subset of music that can be expressed in this format is large enough to be useful for the community of OMR.
In this work we have also presented a concrete implementation of a set of metrics for OMR practitioners with the hopes of bringing together the community to finally speak the same language; a lingua franca thanks to which research can be shared and compared fairly and easily. We also hope this might lower the barrier of entry and bring some new interest on the field.
The work that lies ahead now is bulding a corpus of music that can be employed with this format into a benchmark for CWMN recognition, both in typeset and handwritten domains. For typeset scores, building a corpus in MTN should be as straightforward as compiling a comprehensive set of scores in MusicXML and engraving them. For handwritten music, the million dollar question now lies in how to reliably align MTN to the source material and where to find good transcriptions, which shall be the focus of our next efforts.

### v2

@@ 5.4 Tier 3: Semantic Reconstruction
Time Precision (TP) is defined analogously.
@@
Time Average Shift (TAS) is defined analogously. Signedness is kept in order to identify the direction in which the underlying OMR system tends to move the notes.
@@ 6 The COMREF Dataset
We have developed a dataset (dubbed COMREF from “Common Optical Music Recognition Evaluation Framework”) built on transcriptions of public domain works as a proof of concept of the notation format. In particular, we have used the OpenScore project’s transcriptions of widely known works such as The Art of the Fugue by J.S. Bach or the Planets by Gustav Holst, among others. We have also incorporated the Lieder Corpus Gotham and Jonas (2022) and the String Quartet corpus str (2023). All these scores are engraved from MusicXML files, but we plan on supporting MEI as a conversion source at some point in the future.
@@ 6.1 Images
In summary, the dataset is produced by processing of 894 individual works into images at the measure level (including all staves that belong to it), to produce a total of $435.623$ images. After verification, $461$ images are removed due to noise in the annotations, totalling $435.162$ images after cleanup.
@@
The process through which the dataset was produced is summarised in Figure 4. Scores are engraved through Verovio Pugin (20??), a MEI-based score engraving system which is also compatible with MusicXML, into page-level SVG files. Using the hierarchical structure within the SVG and exploiting the optional identifier information Verovio can be instructed to attach, we employ a splitter script that finds bounding boxes for every measure and engraves them individually. It also finds what measures lie at the beginning of a line in order to properly add this information on the notation. The bounding boxes for all measures are cropped at the next staff vertically and a fixed amount of pixels horizontally.
@@
Figure 5: Some example measures sampled from the dataset. The dataset contains scores of varying types. On the left, an example of a pianoform score with three staves. In the middle, two examples of single-staff scores where the top one is part of a system of staves. On the right, a two-staff pianoform with two staves.
Some qualitative examples of the generated samples can be seen in Figure 5. The dataset purposefully crops images coarsely, incorporating parts of neighboring staves and surrounding distractors. This is in order to better emulate the result of a layout analysis algorithm, which will be inexact. The page-level images from which the measure-level dataset has been produced is also included for use cases that might require it, as well as the Verovio SVG output.
Overall, the resulting dataset is quite challenging because it incorporates a lot of variability, a selection of highly complex scores and lots of distractors.
@@ 6.2 Ground Truth
The dataset has 88 classes which are represented very unevenly. Noteheads, beams, stems and flags are the most common, adding up to 70% of the entire dataset on their own. Of these, 71 are represented in a predefined test partition containing 52884 measures.
It includes music from a wide range of sources, including scores that are monophonic, polyphonic and pianoform. Diverse instrumentations and styles are also covered, with works for large orchestras coexisting with small-scope chamber music or solo works.
6.3 Baseline Experiment
We conducted a simple proof-of-concept experiment on the COMREF dataset to assess the feasibility of the methodology proposed in this paper. For this purpose, we used an off-the-shelf OMR system to produce a transcription of the test partition and we analysed its results.
The OMR system used for this experiment is Audiveris Project (20??), an Open Source page-level system capable of generating a MusicXML output from a single input image. We used this model because it can generate a full score output in MusicXML, it can be run on the full batch of images directly and offers decent results. Other systems in the community or the literature either omit the semantic reconstruction step or cannot be run on the thousands of images of the test set efficiently. Implementing the score reconstruction step for low-level intermediate music representations – e.g. bounding boxes, text representations – is outside the scope of this work. Moreover, few models are designed to address multiple-staff polyphonic music.
The page-level images of the dataset are used as input, since Audiveris requires the information of the clef, key and beat. The output MusicXML is then converted to MTN and a simple matching between predicted and ground truth samples is generated by imposing a top-down reading order given the samples known to be present on each page. If a prediction has more measures per page than the ones in the ground truth, the extra ones are just discarded.
With the setup outlined above, Audiveris predicted 45822 measures from the 52884 present on the ground truth. Out of these, 40622 measures from both sets could be matched together, corresponding to a coverage of 76.9%. The missed predictions are as a result of the engine failing to give an output on certain pages.
Given this setup, the results for Tier 1 are shown in Table 3 and the results for Tiers 2 and 3 are shown in Table 4.
Table 3: Results for Tier 1 computation on the Dataset’s test partition.
Token Precision Recall Counts Prop Token Precision Recall Counts Prop notehead_black 0.946 0.791 193879 0.2992 barline_tok_heavy 0.812 0.913 643 0.0010 stem_down 0.925 0.816 99692 0.1539 dyn_pp 0.901 0.714 597 0.0009 stem_up 0.896 0.756 78848 0.1217 fermata 0.960 0.577 589 0.0009 beam 0.915 0.769 48918 0.0755 tuplet_start 0.239 0.393 563 0.0009 flag 0.922 0.778 28190 0.0435 dyn_ff 0.955 0.813 465 0.0007 notehead_white 0.582 0.715 20330 0.0314 dyn_mf 0.891 0.742 430 0.0007 accidental_sharp 0.941 0.507 19882 0.0307 trill 0.884 0.637 408 0.0006 slur_start 0.852 0.685 17329 0.0267 rest_maxima 0.126 0.648 244 0.0004 slur_stop 0.850 0.675 17320 0.0267 dyn_sfz 0.900 0.117 230 0.0004 accidental_flat 0.907 0.412 15350 0.0237 wavy_line 0.000 0.000 200 0.0003 rest_eighth 0.931 0.742 13821 0.0213 notehead_cue_black 0.000 0.000 185 0.0003 dot 0.904 0.685 10894 0.0168 rest_32nd 0.957 0.817 164 0.0003 staccato 0.637 0.641 9150 0.0141 timesig_common 0.464 0.634 131 0.0002 accidental_natural 0.926 0.738 8537 0.0132 rest_whole 0.082 0.047 106 0.0002 rest_quarter 0.920 0.794 8092 0.0125 repeat_backward 0.653 0.752 105 0.0002 tied_stop 0.822 0.496 7784 0.0120 accidental_double_flat 1.000 0.077 104 0.0002 tied_start 0.812 0.466 7745 0.0120 accidental_double_sharp 1.000 0.646 96 0.0001 clef_G 0.882 0.354 5974 0.0092 clef_oct_F 0.000 0.000 81 0.0001 clef_F 0.554 0.243 3120 0.0048 dyn_fz 0.000 0.000 78 0.0001 rest_long 0.547 0.695 3039 0.0047 dyn_mp 0.846 0.786 70 0.0001 wedge_stop 0.892 0.514 2849 0.0044 clef_oct_G 0.279 0.185 65 0.0001 rest_breve 0.861 0.519 2761 0.0043 repeat_forward 0.295 0.316 57 0.0001 accent 0.802 0.677 2243 0.0035 dyn_sffz 0.000 0.000 47 0.0001 barline_tok_regular 0.787 0.899 2023 0.0031 dyn_ppp 0.000 0.000 44 0.0001 dyn_p 0.771 0.716 1707 0.0026 timesig_cut 0.818 0.878 41 0.0001 dyn_f 0.867 0.698 1669 0.0026 dyn_fff 0.000 0.000 39 0.0001 arpeggiate 0.883 0.062 1573 0.0024 dyn_rf 0.000 0.000 35 0.0001 rest_half 0.908 0.803 1488 0.0023 dyn_fp 0.947 0.581 31 0.0000 wedge_diminuendo 0.874 0.641 1445 0.0022 notehead_breve 0.000 0.000 24 0.0000 wedge_crescendo 0.907 0.710 1433 0.0022 notehead_cue_white 0.000 0.000 11 0.0000 rest_16th 0.936 0.819 1418 0.0022 segno 0.000 0.000 2 0.0000 notehead_grace_black 0.250 0.001 1297 0.0020 dyn_pppppp 0.000 0.000 1 0.0000 clef_C 0.423 0.442 954 0.0015 coda 0.000 0.000 1 0.0000 dyn_sf 0.609 0.671 671 0.0010 turn 0.083 1.000 1 0.0000 tenuto 0.457 0.673 651 0.0010 caesura 0.000 0.000 1 0.0000 Total 0.894 0.733 647965 1.0
Table 4: Results for Tiers 2 and 3. From left to right, Tree Error Rate, Average Time Shift, Average Pitch Shift, Time Precision, Pitch Precision, Staff Precision, False Positive Rate and Missing Note Rate
TER Time Shift Pitch Shift Staff Shift Time Prec. Pitch Prec. Staff Prec. FPR MNR
0.372 -0.096 -0.091 0.022 0.802 0.749 0.963 0.097 0.216
From Tier 1, a few conclusions can be extracted. The first one is that Audiveris can find the structural music elements very reliably (noteheads, stems, beams and flags), all of them with recalls over 75%. This is consistent with the values seen in the missing note rate at roughly 20%, indicating that a majority of the detected noteheads is properly constructed and matched against noteheads in the ground truth. The system is also precise characterising the notes that can be matched, showing a pitch and time shift of roughly $-0.09$. This means that out of every 10 notes, only one is a single step below what it should.
The tree error rate is sensibly high, at $0.372$. This is attributed particularly to the missing elements. Upon inspection of certain samples, the most recurrent error is the omission of certain note groups, causing all of their attached elements to be missed.
The reason for the low scores in attribute elements (clefs, accidentals and time signature elements) is the fact that MusicXML does not encode start-of-line attribute elements because they are layout-dependent. By artificially inserting the tokens corresponding to start-of-sequence elements, we can adjust this to increase the recall but since the matching is sometimes imperfect we also insert a considerable amount of false positives.
Overall, even if the results for this specific tool on the dataset still leave room for improvement, we consider that our proposed format and metric fulfil their design purposes: unique representation of scores and evaluation. Therefore, we consider this simple trial successful.
@@ 7 Conclusions
In this paper we have argued for the implementation of a Common Optical Music Recognition Framework through the instauration of a notation format in which score reconstruction is independent from the recognition methodology. Moreover, the resulting scores can be evaluated fairly an unambiguously. Our proposed reification of this idea is the MTN format, which draws heavy inspiration from existing formats and adapts many ideas already present in the field while also being quite expressive and general. Its main idea is exploiting the expressiveness of graphs while keeping the ordering properties of sequences. Since this method builds upon some of the most widely used abstractions of the community (e.g. symbols as combinations of primitives, time from ordering, etc) it stands as a good candidate for a common endpoint for OMR as a whole. Of course, CWMN is a tremendously complex notation system which has been optimised and streamlined for hundreds of years. Nevertheless, we believe the subset of music that can be expressed in this format is large enough to be useful for the community.
In this work, we have also presented a concrete implementation of a set of metrics for OMR practitioners with the hopes of bringing together the community to speak the same language; a lingua franca thanks to which research can be shared and compared fairly and easily. We provide a simple baseline from which to demonstrate how the evaluation framework works. We also hope this might lower the barrier of entry and bring some new interest in the field.
The work that lies ahead now is building a corpus of music that can be employed with this format into a benchmark for CWMN recognition, both in typeset and handwritten domains. For typeset scores, building a corpus in MTN is as straightforward as compiling a comprehensive set of scores in MusicXML and engraving them. For handwritten music, the main challenge lies in developing techniques to reliably align MTN to the source material andfinding good transcriptions, which shall be the focus of our next efforts.
@@
This work has been partially supported by the Spanish projects PID2021-126808OB-I00 (GRAIL) and CNS2022-135947 (DOLORES). Pau Torras is funded by the Spanish FPU Grant FPU22/00207 and developed part of this work while funded by the AGAUR Joan Oró FI grant 2023 FI-1-00324. The authors acknowledge the support of the Generalitat de Catalunya CERCA Program to CVC’s general activities.